A runner jogs 12km north then turns & runs 16km east in 3 hour.

A) what is his displacement

B) what is his average speed

C) what is his average velovity including the direction


1
Expert's answer
2021-04-07T10:09:29-0400

(A) We can find the displacement of the runner from the Pythagorean theorem:


"d=\sqrt{d_x^2+d_y^2}=\sqrt{(16\ km)^2+(12\ km)^2}=20\ km."

We can find the direction of displacement from the geometry:


"\theta=cos^{-1}(\dfrac{d_x}{d})=36.9^{\circ}\ N\ of\ E."

(B) Average speed of the runner can be found as follows:


"v_{avg}=\dfrac{Total\ distance}{Total\ time}=\dfrac{d_1+d_2}{t},""v_{avg}=\dfrac{12\ km+16\ km}{3\ h}=9.33\ \dfrac{km}{h}."

(C) Average velocity of the runner can be found as follows:


"\vec{v}_{avg}=\dfrac{Total\ displacement}{Total\ time},""\vec{v}_{avg}=\dfrac{20\ km\ [36.9^{\circ}\ N\ of\ E]}{3\ h}=6.67\ \dfrac{m}{s}\ [36.9^{\circ}\ N\ of\ E]."